# Breaking degeneracies in exoplanetary parameters through self-consistent atmosphere-interior modelling

### v2

@@ 9 Conclusions
We present a combined atmosphere and interior model, HADES (Heat Atmosphere Density Evolution Model), that will serve as a powerful tool to give potential structural configurations of Jupiter-like planets. We linked the atmosphere and interior in a self-consistent manner by converging the physical parameters at the interface. We could subsequently reevaluate the intrinsic temperatures corresponding to the internal fluxes of a range of exoplanets. This led us to the conclusion that there is reduced need for ohmic dissipation or other mechanisms that lead to higher intrinsic temperatures to explain the observed radius. We showed that, by using a grid of models, we can use Bayesian inference to derive potential parameters for observed exoplanets. The case of WASP-39 b shows the importance of using spectra to constrain parameters and hence reinforces the importance of atmospheric surveys such as the one anticipated as part of the Ariel mission. The case of 51 Eri b shows the usage a physically coherent model in order to scale the flux with radius correctly even if the fit is of lower quality. 51 Eri b also shows the possibility of inferring formation processes when the systems age is available. Further work is required to complete the model, most notably by adding clouds.

### v3

@@ 9 Conclusions
We present a combined atmosphere and interior model, HADES (Heat Atmosphere Density Evolution Model), which will serve as a powerful tool for proposing potential structural configurations of Jupiter-like planets. We linked the atmosphere and interior in a self-consistent manner by converging the physical parameters at the interface. We subsequently reevaluated the intrinsic temperatures corresponding to the internal fluxes of a range of exoplanets. This led us to the conclusion that there is reduced need for ohmic dissipation or other mechanisms that lead to higher intrinsic temperatures to explain the observed radius. We show that, by using a grid of models, we can use Bayesian inference to derive potential parameters for observed exoplanets. The case of WASP-39 b demonstrates the importance of using spectra to constrain parameters and hence reinforces the importance of atmospheric surveys such as the one anticipated as part of the Ariel mission. The case of 51 Eri b showcases the use of a physically coherent model to correctly scale the flux with radius even if the fit is of lower quality. Our application to 51 Eri b also demonstrates the ability of the interior–atmosphere model to infer formation processes when a system’s age is available. Further work is required to complete the model, most notably by adding clouds.